OpenStudy (anonymous):

If we want to know the velocity that an object is traveling, we must know the

OpenStudy (anonymous):

i'm not sure what you are trying to ask exactly...

OpenStudy (missmob):

to find out the velocity we apply the formula v=d/t mean velocity=displacement divided by/ time taken okay now ill do an numerical for you q1)A cyclist completes half round of a circular track of radius 318 m in 1.5 minutes.Find its speed and velocity? solution: radius of track r=318 m time taken t =1 min 30s =90s (1 min =60s +the remaining 30s) distance covered= pi x radius =3.14x318m = 999m displacement =2r =2x318m =636m speed=distance/time =999m/90s =11.1ms-1 velocity=displacement/time taken =636m/90s =7.07 ms-1 speed=11.1ms-1 velocity= 7.07ms-1